OpenStudy (anonymous):

g(x)= 5x^2 -16x +3 find: g(x+h) - g(x) all divided by h

OpenStudy (luigi0210):

The derivative by definition?

OpenStudy (anonymous):

no you have to plug in g(x) into the equation and I dont know how to do it for the first part

OpenStudy (anonymous):

\[g(x)= 5x^2 -16x +3 \\ g(\heartsuit)= 5\heartsuit^2 -16\heartsuit +3 \\ g(\spadesuit)= 5\spadesuit^2 -16\spadesuit+3 \\ g(\diamondsuit)= 5\diamondsuit^2 -16\diamondsuit +3 \\ g(\clubsuit)= 5\clubsuit^2 -16\clubsuit+3 \\ g(x+h)= 5(x+h)^2 -16(x+h) +3 \]

OpenStudy (anonymous):

on my study guide it says that the answer is 10x +5h -16

OpenStudy (anonymous):

what i wrote is only the first step

OpenStudy (anonymous):

\[g(x+h)= 5(x+h)^2 -16(x+h) +3\] you have to multiply that out, combine like terms

OpenStudy (anonymous):

then subtract \(g(x)\) from it

OpenStudy (anonymous):

yeah sorry i kinda jumped the gun

OpenStudy (anonymous):

then divide everything by \(h\)

OpenStudy (anonymous):

it is a bunch of algebra from here on in

OpenStudy (anonymous):

okay that makes perfect sense, thank you very much.

OpenStudy (anonymous):

yw don't forget to square properly and don't forget to use parentheses when you subtract \(g(x)\)

OpenStudy (anonymous):

5x^2-16x+3 5^x2-15x-3x+3 factor out 5x(x-3)-(x-3) combine (5x-1)(x-3) are ur answers

OpenStudy (anonymous):

lol no

OpenStudy (anonymous):

sorry wrong problem